### v2

@@ 2.1 Detecting trainability prior to training
For the relative entropy, fig. 3 shows the results for the cutoff $\ell^{*}$ and the accuracy of neural networks for two different cases: a fixed-width ($N_{\ell<L-1}=784$, $N_{L-1}=400$) network for MNIST in (a), and a monotonically shrinking-width network for CIFAR10 in (b). For the latter, the width is $O(10)$ for the last eight layers (excluding the output layer with $N_{L}=10$), at which point the analysis based on the central limit theorem poole2016exponential ; schoenholz2017deep no longer applies. As shown in fig. 3, the cutoff $\ell^{*}$ shown in white - qualitatively predicts the empirical correlation length as measured by the accuracy of the trained network. The exact cutoff value depends on the threshold for numerical fluctuations; we have found that a threshold value of around $\eta=0.005$ provides the best quantitative match, though the ordered phase (low $\sigma_{w}^{2}$) appears less sensitive to this choice than the chaotic phase (high $\sigma_{w}^{2}$). This is unsurprising in light of the size of thermal fluctuations in either phase, cf. panel (f) of fig. 2. Notably, the location of the critical point – determined in the reconstruction entropy method as the point at which $\ell^{*}=L$ – is relatively robust to the choice of threshold value, and exhibits excellent agreement with the critical regime determined by actually training the DNN, even in the finite-width regime (note the shift in the peak relative to the central limit theorem prediction in panel (b), potentially due to finite-width effects).
@@
Figure 3: Predicting trainability with reconstruction relative entropy. The figure shows trainability on MNIST (panel (a), fixed width) and CIFAR10 (panel (b), tapering) along a fixed $\sigma_{b}^{2}=0.05$ slice through the 2d phase space $(\sigma_{w}^{2},\sigma_{b}^{2}$), cf. schoenholz2017deep . The colour corresponds to the final accuracy obtained by the networks after 200 epochs of training. Overlaid is the reconstruction relative entropy cutoff $\ell^{*}$ for a range of threshold values (white curve), as well as the $N\to\infty$ central limit theorem prediction (black curves). The relative entropy cutoff accurately predicts the location of the critical point, providing a proxy for the correlation length in the system. We emphasize that the reconstruction entropy curves are obtained prior to training the DNN, i.e., for untrained networks, and are thus orders of magnitude cheaper to generate. Note that the sharp drop-off in the second panel is consistent with fig. 7 of schoenholz2017deep , and can be smoothed with significantly longer training times, cf. fig. 4.
@@
Figure 4: Predicting trainability with reconstruction differential entropy. The figure shows the measurement of differential entropy given by (4) for a wide neural network 200 layers deep, trained on MNIST. The entropy is calculated via the reconstruction method prior to any training. As shown in the color scheme to the right, red corresponds to large information loss (i.e., untrainable), while blue corresponds to more information being preserved (i.e., more trainable), as quantified by the differential entropy. In the plot, the ordered phase is clearly visible (left side) due to the rapid decay of correlations. This rapid information loss causes a clear delineation between the trainable and untrainable regimes. In the chaotic regime (right side), information is lost much more gradually, such that sufficiently long training will eventually allow deeper networks to train successfully, albeit with potentially significant costs (c.f. (schoenholz2017deep, , fig. 7)). Around the critical point of $\sigma_{w}^{2}=1.76$, information is preserved even for deeper layers (blue cone). These predictions, obtained prior to training, are in agreement with the empirical observations for trained networks shown in schoenholz2017deep .
@@
In fig. 4, the differential entropy is shown for a wide neural network with 200 layers as function of the weight variance. The network is untrained and the differential entropy is calculated for the MNIST data. The numerical stability of the differential entropy is indeed significantly improved as compared to the relative entropy shown in fig. 3. In fig. 4, the three different phases can easily be identified: On the left hand side, the network is in the ordered phase and information is lost within a few ($O(1)$) layers. This is reflected in a large negative differential entropy. At the critical point, the differential entropy remains almost unchanged over the depth of the network. In the chaotic regime, the differential entropy decreases with increasing depth. However, the information loss is slower than in the ordered regime. This can be seen as the differential entropy in the ordered regime is around $-18$ for the deeper networks, whereas for the chaotic phase similar deep networks cause a differential entropy of around $-13$. Moreover, the behaviour is less stable and more fluctuations are observed. The image shown again highlights the relation between information content and trainability: Close to criticality the information loss over depth is limited and the network is thus able to learn. For the other regions the information is lost at some depth, reducing the trainability. As the evaluation of trainability depends on a number of maximal epochs to train, the exact information content for training a network is unknown. However, the more information reaches a given depth of the network, the faster the network will be able to learn and the higher its accuracy will be.
@@ 2.2 Reduction in overall training time
In practice, the choice of hyperparameters such as the learning rate also play an important role in trainability (hence the widespread use of optimizers such as ADAM ADAM ). In general, the brute-force method for determine the optimum initial conditions is to perform a computationally costly grid search over all possible combinations. One example of optimization algorithms is optuna [^10] See optuna_2019 and https://optuna.readthedocs.io/en/stable/. The associated time complexity is a product of the number of values of each (hyper)parameter. As a simplified example, if we limit ourselves to $M_{w}$ values of the weight variance $\sigma_{w}^{2}$ and $M_{r}$ values of the learning rate $r$, then the time complexity for this search is $O(M_{w}M_{r})$. From the phase space perspective taken here however, it is more efficient to first determine the trainable regime in parameter space, and only then select the optimum hyperparameters, cf. footnote 1. This reduces the time complexity to $O(M_{w}+M_{r})$. [^11] More generally, we have $O\left(M_{w}M_{b}\prod_{i}M_{i}\right)\to O\left(M_{w}M_{b}+\prod_{i}M_{i}\right)$ where $M_{b}$ is the number of bias variances and $M_{i}$ the number of the $i^{\mathrm{th}}$ hyperparameter over which we scan. In other words, we split the grid search into a search over phase space $(\sigma_{w}^{2},\sigma_{b}^{2})$ followed by a search over hyperparameters. At each point in the former, we enjoy the speedup provided by the RE method for determining whether a network at that point is trainable for some choice of the latter. [^12] That is, initialization in the critical regime (more generally, $L\leq\ell^{*}$) is a necessary but not sufficient condition for trainability in the sense that one must still select hyperparameters, but – assuming one is below the line of uniformity identified in Bukva:2023ksv – it is a necessary and sufficient condition in the sense that some choice of hyperparameters will work. In fact, we expect the time complexity of locating the critical value of $\sigma_{w}^{2}$ itself to be nearer $O(\ln M_{w})$, since the form of the cutoff curves in fig. 3 is amenable to a binary search, modulo fluctuations.
@@ 2.3 Opening the black box: retracing the neural network’s decisions
Figure 5: Visualizations of the network’s decision process. (a) Reconstruction of the sequence of layers of a shrinking neural network with 10 hidden layers, initialized with $\sigma_{w}=1.65$ and trained for 20 epochs. The variance of $\sigma_{w}=1.65$ is chosen close to the critical point and thus in the well-trainable regime. The input is drawn from the MNIST dataset. The input image is labelled as a 5. As can be seen, the features of the digit change over the depth of the network. Most prominently, the left upper edge at the horizontal bar of the 5 becomes more sharp. (b) Same as (a), except that here the network predicts a 6, which is shown in the lower row of the panel. The network adds a smooth connection between two sections of the 5, causing the internal representation to resemble a 6. (c) Reconstruction of the optimal digits as represented by the network.
@@
Moreover, we may use our method to reconstruct the images that the network considers perfect examples of the corresponding label, thereby gaining insight into how this information is represented internally. To this end, we create an output vector with maximal activation on the label chosen as example. Fig. 5, panel (c), shows the optimal digits for the given network.
@@ 3 Discussion
An obvious technical improvement would be to train the reconstruction layers in parallel. The gains from parallelization are highly hardware dependend and we have not explored this here. For networks with $O(10^{2})$ layers, this could potentially lead to a further speedup of two additional orders of magnitude. Additionally, we have demonstrated the effectiveness of this reconstruction method only for fully-connected feedforward networks (MLPs). A further taks is to expand this method to more complex architectures, such as convolutional neural networks, recurrent neural networks, and transformers.
@@ A.2 Determining the cutoff
However, the saturation is not perfect, since the entropy continues to exhibit small fluctuations about this asymptotic value over all remaining layers. Therefore, to determine the cutoff numerically, we must choose a tolerance $\eta$, and proceed as follows: we set $\ell^{*}=L$ by default, and store the associated reconstruction entropy (RE). We then sequentially decrease the layer index $\ell_{i}$ until the difference between the RE at layer $L$ and that at layer $\ell_{i}$ exceeds the threshold $\eta$. At this point, we set $\ell^{*}=\ell_{i}+1$ (so that if the threshold is immediately exceeded, we recover $\ell^{*}=L$).
@@
Acknowledgements. We are grateful to Haye Hinrichsen, Luca Kohlhepp, Dominik Neuenfeld and Martin Rackl for useful discussions. JE and YT acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, project-id 390858490). The work of JE and YT is also supported by DFG grant ER 301/8-1.
Author contributions. The project was jointly initiated by all three authors. YT performed all calculations involved. JE and RJ provided supervision and guidance. All three authors contributed to writing the manuscript.

### v3

@@ 2.1 Detecting trainability prior to training
For the relative entropy, fig. 3 shows the results for the cutoff $\ell^{*}$ and the accuracy of neural networks for a fixed-width ($N_{\ell<L-1}=784$, $N_{L-1}=400$) network for MNIST. As shown in fig. 3, the cutoff $\ell^{*}$ shown in white - qualitatively predicts the empirical correlation length as measured by the accuracy of the trained network. The exact cutoff value depends on the threshold for numerical fluctuations (see Methods below); we have found that a threshold value of around $\eta=0.005$ provides the best quantitative match, though the ordered phase (low $\sigma_{w}^{2}$) appears less sensitive to this choice than the chaotic phase (high $\sigma_{w}^{2}$). This is unsurprising in light of the size of thermal fluctuations in either phase, cf. panel (f) of fig. 2. Notably, the location of the critical point – determined in the reconstruction entropy method as the point at which $\ell^{*}=L$ – is relatively robust to the choice of threshold value, and exhibits excellent agreement with the critical regime determined by actually training the DNN.
@@
Figure 3: Predicting trainability with reconstruction relative entropy. The figure shows trainability on MNIST along a fixed $\sigma_{b}^{2}=0.05$ slice through the 2d phase space $(\sigma_{w}^{2},\sigma_{b}^{2}$), cf. schoenholz2017deep . The colour corresponds to the final accuracy obtained by the networks after 200 epochs of training. Overlaid is the reconstruction relative entropy cutoff $\ell^{*}$ for a threshold value of $\eta=0.005$ (white curve), as well as the $N\to\infty$ central limit theorem prediction (black curves). The relative entropy cutoff accurately predicts the location of the critical point, providing a proxy for the correlation length in the system. We emphasize that the reconstruction entropy curves are obtained prior to training the DNN, i.e., for untrained networks, and are thus orders of magnitude cheaper to generate.
@@
Figure 4: Predicting trainability with reconstruction differential entropy. The figure shows the measurement of differential entropy given by (4) for a wide neural network 200 layers deep, trained on MNIST. The entropy is calculated via the reconstruction method prior to any training. As shown in the color scheme to the right, red corresponds to large information loss (i.e., untrainable), while blue corresponds to more information being preserved (i.e., more trainable), as quantified by the differential entropy. The ordered phase is clearly visible (left side) due to the rapid decay of correlations. This rapid information loss causes a clear delineation between the trainable and untrainable regimes. In the chaotic regime (right side), information is lost much more gradually, such that sufficiently long training will eventually allow deeper networks to train successfully, albeit with potentially significant costs (c.f. (schoenholz2017deep, , fig. 7)). Around the critical point of $\sigma_{w}^{2}=1.76$, information is preserved even for deeper layers (blue cone). These predictions, obtained prior to training, are in agreement with the empirical observations for trained networks shown in schoenholz2017deep .
Figure 5: Predicting trainability for a ResNet with reconstruction differential entropy. The figure shows the predicted trainability with reconstruction differential entropy for a ResNet trained on FashionMNIST. The bias variance is set to $\sigma_{b}^{2}=5e^{-2}$. We observe a similar phase behaviour as in fig. 4, though the edges of the trainable region are less sharp. Particularly in the chaotic phase (towards the right), the large fluctuations seem to indicate the role of residual connections in preserving relevant information despite the otherwise limited signal propagation. Nonetheless, one clearly sees an optimum critical region at approximately $\sigma_{w}^{2}\sim 1.3$
@@
In fig. 4, the differential entropy is shown for a wide neural network with 200 layers as function of the weight variance. The network is untrained and the differential entropy is calculated for the MNIST data. The numerical stability of the differential entropy is indeed significantly improved as compared to the relative entropy shown in fig. 3. In fig. 4, the three different phases can easily be identified: On the left hand side, the network is in the ordered phase and information is lost within a few ($O(1)$) layers. This is reflected in a large negative differential entropy. At the critical point, the differential entropy remains almost unchanged over the depth of the network. In the chaotic regime, the differential entropy decreases with increasing depth. However, the information loss is slower than in the ordered regime. This can be seen as the differential entropy in the ordered regime is around $-18$ for the deeper networks, whereas for the chaotic phase similar deep networks cause a differential entropy of around $-13$. Moreover, the behaviour is less stable and more fluctuations are observed. The image shown again highlights the relation between information content and trainability: close to criticality the information loss over depth is limited and the network is thus able to learn. For the other regions the information is lost at some depth, reducing the trainability. As the evaluation of trainability depends on a number of maximal epochs to train, the exact information content for training a network is unknown. However, the more information reaches a given depth of the network, the faster the network will be able to learn and the higher its accuracy will be. Figure 5, shows similar results for a ResNet trained on FashionMNIST. The bias variance is set to $\sigma_{b}^{2}=2e^{-5}$, similar to the CNN in xiao2018dynamical as its residual blocks consist of convolution layers. The ResNet consists of 50 residual blocks. Each block consists of 2 convolutional layers (cf. he2015deepresiduallearningimage ). Ignoring the skip connections, the ResNets are thus similar to 100 layer deep convolutional networks.
@@ 2.2 Reduction in overall training time
In practice, the choice of hyperparameters such as the learning rate also play an important role in trainability (hence the widespread use of optimizers such as ADAM ADAM ). In general, the brute-force method for determine the optimum initial conditions is to perform a computationally costly grid search over all possible combinations. One example of optimization algorithms is optuna [^12] See optuna_2019 and https://optuna.readthedocs.io/en/stable/. The associated time complexity is a product of the number of values of each (hyper)parameter. As a simplified example, if we limit ourselves to $M_{w}$ values of the weight variance $\sigma_{w}^{2}$ and $M_{r}$ values of the learning rate $r$, then the time complexity for this search is $O(M_{w}M_{r})$. From the phase space perspective taken here however, it is more efficient to first determine the trainable regime in parameter space, and only then select the optimum hyperparameters, cf. footnote 2. This reduces the time complexity to $O(M_{w}+M_{r})$. [^13] More generally, we have $O\left(M_{w}M_{b}\prod_{i}M_{i}\right)\to O\left(M_{w}M_{b}+\prod_{i}M_{i}\right)$ where $M_{b}$ is the number of bias variances and $M_{i}$ the number of the $i^{\mathrm{th}}$ hyperparameter over which we scan. In other words, we split the grid search into a search over phase space $(\sigma_{w}^{2},\sigma_{b}^{2})$ followed by a search over hyperparameters. At each point in the former, we enjoy the speedup provided by the RE method for determining whether a network at that point is trainable for some choice of the latter. [^14] That is, initialization in the critical regime (more generally, $L\leq\ell^{*}$) is a necessary but not sufficient condition for trainability in the sense that one must still select hyperparameters, but – assuming one is below the line of uniformity identified in Bukva:2023ksv – it is a necessary and sufficient condition in the sense that some choice of hyperparameters will work. In fact, we expect the time complexity of locating the critical value of $\sigma_{w}^{2}$ itself to be nearer $O(\ln M_{w})$, since the form of the cutoff curves in fig. 3 is amenable to a binary search, modulo fluctuations.
@@ 2.3 Opening the black box: retracing the neural network’s decisions
Figure 6: Visualizations of the network’s decision process. (a) Reconstruction of the sequence of layers of a shrinking neural network with 10 hidden layers, initialized with $\sigma_{w}=1.65$ and trained for 20 epochs. The variance of $\sigma_{w}=1.65$ is chosen close to the critical point and thus in the well-trainable regime. The input is drawn from the MNIST dataset. The input image is labeled as a 5. As can be seen, the features of the digit change over the depth of the network. Most prominently, the left upper edge at the horizontal bar of the 5 becomes more sharp. (b) Same as (a), except that here the network predicts a 6, which is shown in the lower row of the panel. The network adds a smooth connection between two sections of the 5, causing the internal representation to resemble a 6. (c) Reconstruction of the optimal digits as represented by the network.
@@
Moreover, we may use our method to reconstruct the images that the network considers perfect examples of the corresponding label, thereby gaining insight into how this information is represented internally. To this end, we create an output vector with maximal activation on the label chosen as an example. Fig. 6, panel (c), shows the optimal digits for the given network.
As a more complex proof of concept, we apply the reconstructions to the convolutional network vgg16 simonyan2015deepconvolutionalnetworkslargescale . The trained weights for this network are taken from the torch library torchlib . We train the cascade networks by $30,000$ examples form the ImageNet dataset 5206848 for a single epoch each. We use the test image of Selvaraju_2019 , displaying a cat and a dog in front of a window as an example. Using these cascade networks, we are able to reconstruct the information propagating to any given layer, as well as the information within individual channels. We use the activations in all channels of a given layer of the CNN for the reconstruction (cf. fig. 7 b). Alternatively, we may suppress the information in different channels and focus on a single one instead (cf. fig. 7 a). The reconstruction method allows us to determine which properties – e.g., degree of brightness, presence of edges or animals – are stored in different channels. For example, in fig. 7 a), a selection of reconstructions of different channels in different layers of the CNN are shown: in the top right panel, titled “layer 3 channel 26” the brightness is memorized by the network. This includes the bright spot on the floor from the window and the left side of the dog. In the lower left panel, titled “layer 6 channel 3”, the channel focuses on edges. In the lower right panel, titled “layer 10 channel 112”, bright spots mostly belonging to the cat and dog are focused on. Moreover, in fig. 7 b), the full reconstruction, i.e. the reconstruction using all channels after all convolution and pooling layers, is shown. Obviously, the original image is now blurred. That is expected as vgg16 involves multiple Maxpooling layers, causing a loss of local information and the network in general is expected to decrease the information with increasing depth. In this case, vgg16’s most probable prediction for the test image is the class with label 243, which predicts a “bull mastiff” in the image.
Figure 7: Panel (a) shows selected reconstructions of different channels of different layers for the input image and the vgg16 convolutional network simonyan2015deepconvolutionalnetworkslargescale . The original image is taken from Selvaraju_2019 . For each layer we present a single example of channel, which shows a particular characteristic of the image (e.g., brightness and edges). Different channels in the layers yield different features of the image. Panel (b) shows a reconstruction using all channels, starting at the last pooling layer at a depth of 17.
@@ 3 Discussion
An obvious technical improvement would be to train the reconstruction layers in parallel. The gains from parallelization are highly hardware-dependent and we have not explored this here. For networks with $O(10^{2})$ layers, this could potentially lead to a further speedup of two additional orders of magnitude. Additionally, we have demonstrated the effectiveness of this reconstruction method only for fully-connected feedforward networks (MLPs). A further task is to expand this method to more complex architectures, such as convolutional neural networks, recurrent neural networks, and transformers.
@@ A.2 Determining the cutoff
However, the saturation is not perfect, since the entropy continues to exhibit small fluctuations about this asymptotic value over all remaining layers. Therefore, to determine the cutoff numerically, we must choose a tolerance or threshold $\eta$, and proceed as follows: we set $\ell^{*}=L$ by default, and store the associated reconstruction entropy (RE). We then sequentially decrease the layer index $\ell_{i}$ until the difference between the RE at layer $L$ and that at layer $\ell_{i}$ exceeds the threshold $\eta$. At this point, we set $\ell^{*}=\ell_{i}+1$ (so that if the threshold is immediately exceeded, we recover $\ell^{*}=L$). As described in the main text, we have found that a value of $\eta=0.005$ works well in practice. If the value chosen for $\eta$ is too small, it will be immediately exceeded and one will have $\eta^{*}=L$; conversely, if the value is too large, the fluctuations will dominate over the curve. Thus, choosing the optimum value of $\eta$ is tantamount to selecting a suitable signal-to-noise ratio for the experiment.
@@
A.3 Entropy of white noise
To further illustrate that the information propagation is primarily a question of the location of the network in phase space rather than a property of the dataset itself, in fig. 8 we show the predicted trainability with reconstruction differential entropy for a DNN fed white noise. Again, we do not need to train the feedforward network: the point is that the cascades can still be trained to reconstruct the input at each layer, and that the associated reconstruction entropy computed with our method shows excellent agreement with fig. 4, in which the dataset was MNIST.
Figure 8: Predicting trainability for a wide neural network fed white noise with reconstruction differential entropy. Here we show the same network architecture as in fig. 4, but instead of taking a structured dataset like MNIST, we feed the network white noise (again, no training of the feedforward network is performed). Despite this, the reconstruction differential entropy computed via the cascade layers is still able to distinguish the trainable regime.
A.4 Entropy of a CNN
Fig. 9 shows the predicted trainability with differential entropy for a CNN trained on MNIST. For a better visualization the values for the differential entropy in the ordered phase are clipped at -10 to increase contrast.
Figure 9: Predicting trainability for a CNN with reconstruction differential entropy. Again we observe similar phase behaviour as in fig. 4, with a sharp delineation at the ordered phase boundary (left side), and a broad region of sub-optimal trainability in the chaotic phase right side). The optimum critical region is still visible as the densest peak near the ordered phase boundary.
A.5 Pixel activations over images
The plots in Fig. 10 substantiate the theoretical argument given above (4) that after reconstruction, the distribution of the activation of given pixel across images is approximately Gaussian. For a given pixel, we plot the distribution of the activation over a random batch of MNIST images after reconstruction. The histograms are shown for various layers in a DNN initalized at an arbitrary point in phase space.
Figure 10: Histograms for the values of pixels 0 (left column), 391 (middle column), and 783 (right column) at layer 1 (top row), 10 (middle row), and 99 (bottom row) for a network initialized with variance $\sigma_{w}^{2}=3.4$. This corresponds to a network in the chaotic phase, but these plots are qualitatively unchanged at other points in phase space (e.g., in the ordered or critical regime).
Acknowledgments. We are grateful to Haye Hinrichsen, Luca Kohlhepp, Dominik Neuenfeld and Martin Rackl for useful discussions. JE and YT acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, project-id 390858490). The work of JE and YT is also supported by DFG grant ER 301/8-1.
Author contributions. The project was jointly initiated by all three authors. YT performed all coding involved. RJ and JE provided supervision and guidance. All three authors contributed to writing the manuscript.